# Minimal Thermodynamic Cost of Computing with Circuits

### v2

@@ Abstract
All digital devices have components that implement Boolean functions, mapping that component’s input to its output. However, any fixed Boolean function can be implemented by an infinite number of circuits, all of which vary in their resource costs. This has given rise to the field of circuit complexity theory, which studies the minimal resource cost to implement a given Boolean function with any circuit. Traditionally, circuit complexity theory has focused on the resource costs of a circuit’s size (its number of gates) and depth (the longest path length from the circuit’s input to its output). In this paper, we extend circuit complexity theory by investigating the minimal thermodynamic cost of a circuit’s operation. We do this by using the mismatch cost (MMC) of a given circuit that is run multiple times in a row to calculate a lower bound on the entropy production (EP) incurred in each such run. Specifically, we discuss conditions under which MMC of a circuit is proportional to the size of a circuit, and conditions for them to diverge. We also use our results to compare the thermodynamic costs of different circuit families implementing the same family of Boolean functions. In addition we analyze how heterogeneity in the underlying physical processes implementing the gates in a circuit influences the minimal thermodynamic cost of the overall circuit. These and other results of ours lay the foundation for extending circuit complexity theory to include mismatch cost as a resource cost.
@@ I Introduction
Computational complexity theory explores the relative difficulty of computing functions within formal models of computation. In his foundational work turing1936computable , Alan Turing introduced the concept of Turing machines and demonstrated that functions can be classified as either computable or non-computable. However, even within the class of computable functions, some are inherently harder to compute than others rabin1960degree ; cook2007overview . The fundamental question arises: what does it mean for one function to be more difficult to compute than another?
@@
A given function can often be computed using multiple algorithms, each incurring different resource costs depending on the model of computation. Traditionally, computational complexity theory has focused on quantifying resource usage such as the number of computational steps or memory requirements. Early discussions on complexity measures also considered physical energy expenditure and thermodynamic work as potential complexity metrics cobham1965intrinsic .
@@
Landauer and Bennett were among the first to explore the thermodynamic cost of computation using equilibrium statistical mechanics. They famously argued that erasing a single bit of information necessarily generates at least $k_{B}\ln 2$ of heat, where $k_{B}$ is Boltzmann’s constant and $T$ is the temperature of the surrounding heat bath landauer1961irreversibility ; bennett1982thermodynamics . However, this early work modeled computation—which is inherently a non-equilibrium process—within the framework of equilibrium thermodynamics. As a result, the analysis was semi-formal and limited in scope, leaving many fundamental questions unanswered. At the time, the theoretical tools needed to rigorously analyze far-from-equilibrium systems were not yet available, and the thermodynamics of computation remained an open and underdeveloped area.
@@
Fortunately, recent advances in non-equilibrium thermodynamics, particularly within the framework of stochastic thermodynamics, have made it possible to rigorously define and analyze thermodynamic quantities such as work, entropy, and heat dissipation in systems driven far from equilibrium van2015ensemble ; seifert2012stochastic . These tools are now being applied to fundamental questions about the thermodynamic costs of computation and communication processes van2015ensemble ; seifert2012stochastic ; wolpert2019stochastic ; wolpert2024thermodynamics ; wolpert2024stochastic . In particular, stochastic thermodynamics allows us to analyze the entropy production in a process, i.e., the unavoidably wasted energy incurred by running that process.
@@
Boolean circuits provide an alternative model of computation, characterizing the complexity of a function in terms of the size and depth of the circuit implementing it. The size of a circuit is the total number of gates, while the depth corresponds to the longest path from an input to an output, representing the time required for computation arora2009computational ; vollmer1999introduction . Just as there exist multiple algorithms for a given problem, there are infinitely many Boolean circuits that can implement a given Boolean function. The task of identifying the most efficient circuit among those infinitely many circuits is both theoretically difficult and practically important. Circuit complexity theory provides a framework for comparing and optimizing circuits based on size and depth, but the associated thermodynamic cost of running a circuit as a key measure in optimization has not been given considerable attention.
@@
As processors shrink toward their theoretical efficiency limits—and as computational demands surge, especially due to large-scale data processing—understanding the energy cost of computation is becoming increasingly important. Incorporating thermodynamic considerations into circuit complexity theory offers a principled framework to quantify minimal heat dissipation. This perspective can refine existing complexity measures and inform the design of circuits that are both logically and thermodynamically efficient. Moreover, since circuits determine whether a function is efficiently parallelizable or inherently sequential, thermodynamic analysis can illuminate the energy efficiency of parallel computation.
A central concept in this analysis is the mismatch cost (MMC) kolchinsky2021dependence ; kolchinsky2017dependence , which is a component of the entropy production arising in physical computational processes.
@@
Sec. IV builds on this earlier analysis to introduce mismatch cost complexity as a new circuit complexity measure, analogous to size and depth complexity, representing the minimal energetic cost of implementing a Boolean function with any circuit. We derive results and theorems that link MMC complexity to size complexity, showing that under certain conditions, the upper bound on MMC grows linearly with the size complexity of a circuit family. Complementing this result, we identify conditions where MMC scales differently from size complexity. In Sec.V, we discuss the energetic aspects of circuit optimization by comparing the MMC associated with circuit families that implement a given Boolean function. Specifically, we study two families that compute the ADD function—one with depth complexity $\log(n)$ and another with $n$. We also investigate the impact of heterogeneity in the physical properties of gates on the associated MMC. Finally, Sec. VI concludes the paper by discussing possible directions for future research.
@@ II.1 Circuit complexity theory
A circuit is built from basic logic gates that are interconnected to process inputs and produce outputs. A basis is a set $\mathcal{B}$ of Boolean functions that define the allowable gates in a circuit. For example, a basis containing AND and NOT gates is considered a universal basis because any Boolean function can be implemented using only these gates. Formally, a Boolean circuit is represented as a directed acyclic graph (DAG), denoted by $(V,E,\mathcal{B},\mathcal{X})$. Here, $V$ is the set of nodes, where each node represents a logic gate selected from a finite basis $\mathcal{B}$. The edges in $E$ define the directed connections between gates, determining how information flows through the circuit. The set $\mathcal{X}$ represents the state space of the circuit (see below). The direction of edges reflects dependencies between gates, enabling a specific ordering of nodes in $V$. This ordering, called a topological ordering, ensures that for every directed edge from node $\mu$ to node $\nu$, node $\mu$ appears before node $\nu$ in the sequence ($\mu<\nu$).
A node $\mu$ has incoming edges from nodes known as its parent nodes, denoted as $\mathrm{pa}(\mu)$, and outgoing edges to nodes known as its children nodes, denoted as $\mathrm{ch}(\mu)$. Nodes that have no incoming edges, called input nodes or root nodes, serve as the circuit’s inputs. We denote the set of input nodes as $V_{in}\subset V$ and the set of all non-input nodes as $V_{nin}=V\setminus V_{in}$.
Each gate $\mu$ in the circuit has an associated state space, denoted by $X_{\mu}$, and its state is represented by $\boldsymbol{x}_{\mu}\in X_{\mu}$. For binary gates, the state space is $X_{\mu}=\{0,1\}$. The collective states of all gates together define the joint state of the circuit. The joint state space of the circuit is given by
@@
We use $\boldsymbol{x}_{\mathrm{pa}(\mu)}$ and $\boldsymbol{x}_{\mathrm{ch}(\mu)}$ to denote the joint state of the parent and children gates of a gate $\mu$, respectively. Similarly, we define $\boldsymbol{x}_{in}$ and $\boldsymbol{x}_{nin}$ to represent the joint states of all input nodes and all non-input nodes, respectively.
@@
Circuits differ fundamentally from other models of computation. In models like Turing machines, a single machine accepts input of of any arbitrary length. In contrast, circuits require a distinct specification for each input length. This means that to compute functions belonging to the same family but with different input sizes, one must design distinct circuits for each case. Because the model does not provide a uniform mechanism for handling all input lengths, it is classified as a non-uniform model of computation arora2009computational ; vollmer1999introduction . As a result, computation in this model is described using a circuit family, denoted as $\{C_{n}\}_{n\in\mathbb{N}}$, where each circuit $C_{n}$ corresponds to inputs of length $n$. A circuit family $\mathcal{C}~{}$ is a sequence $\{C_{n}\}_{n\in\mathbb{N}}$ of Boolean circuits, where $C_{n}$ is the circuit with $n$ inputs.
@@
Analogously, let $d:\mathbb{N}\to\mathbb{N}$. $0pt{d}$ is the class of functions $f$ for which there exist a circuit family $\mathcal{C}~{}$ of depth complexity $d$.
The growth of a circuit’s size or depth with respect to the input length of a function provides a crucial measure of the function’s computational complexity. In his seminal work, Shannon proved that most Boolean functions require circuits of exponential size, specifically showing that there exist functions that cannot be computed by circuits smaller than $2^{n}/10n$. In other words, the vast majority of functions—often referred to as "hard" functions—such that $f\notin\mathbf{SIZE}(2^{n}/10n)$ and thus require exponentially large circuits to compute shannon1949synthesis . However, despite this result, explicitly identifying such a function has remained challenging, as most functions of practical interest can still be computed by circuits of reasonable, albeit large, size.
An interesting class of functions emerges when restricted to "small" circuits, i.e., circuit families whose size does not grow faster than polynomial with input length. $\mathbf{P_{/poly}}$ is the class of functions for which there exist a $\mathrm{poly}(n)$-size circuit family,
@@
A function has an efficiently parallel algorithm if and only if it is in $\mathbf{NC}$ arora2009computational ; vollmer1999introduction .
The previous example of ADD is in $\mathbf{NC}$. However, just like any Boolean function, ADD has many circuit families that can implement it. For example, ADD can be implemented using two well-known circuit families: the ripple-carry adder (RCA) and the carry look-ahead adder (CLA), each offering distinct trade-offs in size and depth complexity vollmer1999introduction .
The RCA follows the simple "carry-over" method taught in high school arithmetic, where each bit addition depends on the carry from the previous bit. As a result, for an $n$-bit input, the depth of the circuit grows linearly with $n$. Additionally, increasing the input size by one bit requires adding one additional full-adder circuit to the chain. This means that the overall circuit size also grows linearly with $n$. Its simplicity makes it easy to scale. However, since each stage of bit addition must wait for the carry from the previous stage before proceeding, RCA becomes significantly slower as the input size increases.
In contrast, CLA uses a more sophisticated approach by predicting carry bits in advance rather than waiting for them to propagate sequentially. This significantly reduces the depth of the circuit, which grows proportionally to the logarithm of the input size. As a result, computation time decreases drastically for large inputs. However, the parallel computation and complex wiring cause the total number of gates in the circuit to grow at a rate proportional to $\log n$. Consequently, as the input size increases, the CLA circuit becomes increasingly intricate.
@@ II.2 Stochastic Thermodynamics and Mismatch Cost
Real-world computational systems operate far from equilibrium, with multiple interdependent components evolving rapidly. Traditional equilibrium statistical mechanics, which applies only to large macroscopic systems that either evolve quasi-statically or remain static, is inadequate for analyzing the energetic costs of computation wolpert2024stochastic . Mesoscopic systems far from equilibrium that exhibit stochastic fluctuations can be analyzed within the framework of stochastic thermodynamics. The key assumption in this framework is that any degrees of freedom not explicitly described by the system’s dynamics—such as internal states or the reservoirs—remain in equilibrium. This assumption allows one to relate thermodynamic quantities to system dynamics through the principle of local detailed balance seifert2012stochastic ; seifert2018stochastic ; esposito2010entropy ; esposito2012stochastic .
One of the central quantities in stochastic thermodynamics is entropy production, which quantifies the irreversible heat dissipation in a process. If the probability distribution over the states of a physical system during its evolution changes from $p_{0}$ to $p_{\tau}$, the total EP associated with an initial distribution $p_{0}$ can be expressed as van2015ensemble ; kolchinsky2017dependence ,
@@
Advances in stochastic thermodynamics have led to the discovery of various contributions to total EP based on key properties of a system’s dynamics. Notably, thermodynamic uncertainty relations (TURs) establish a contribution to EP that arises from he precision of currents involved in the system’s evolution barato2015thermodynamic ; gingrich2016dissipation ; horowitz2020thermodynamic . Another significant class of results, known as speed limit theorems (SLTs), provide a lower bound on EP in terms of the average number of state transitions during the system’s evolution shiraishi2018speed ; vo2020unified . In the context of computation, these results depend on the nitty-gritty details of how the map $G$ is implemented—for instance, SLTs require knowledge of the average number of transitions in the underlying dynamics, while TURs depend on the precision of currents involved in the process.
Another key contribution to EP is the mismatch cost. Consider the initial distribution that minimizes EP in Eq. 1, denoted as $q_{0}$, which we refer to as the prior distribution of the process. When the process starts from a different initial distribution $p_{0}$, the resulting EP, $\sigma(p_{0})$, can be decomposed as kolchinsky2017dependence ; kolchinsky2021dependence ,
@@
where $D(p_{0}||q_{0})$ is the Kullback-Leibler (KL) divergence between $p_{0}$ and $q_{0}$, and $\sigma_{\text{res}}$, known as the residual EP, represents the minimum EP of the process, defined as $\sigma_{\text{res}}:=\sigma(q_{0})$. The drop in KL-divergence $D(p_{0}||q_{0})-D(Gp_{0}||Gq_{0})$ is called the mismatch cost. The MMC is always non-negative, a consequence of the data processing inequality for KL-divergence. Unlike the residual EP $\sigma_{\text{res}}$, which depends on the fine details of the physical implementation, MMC depends only on the computational map $G$, the initial distribution $p_{0}$, and the prior distribution $q_{0}$. Additionally, $\sigma_{\text{res}}$ is non-negative due to the second law. In certain dynamical situations, the MMC provides a strictly positive contribution to EP. Consider a process where the computational map $G$ is repeatedly applied to the state space. Since the underlying physical process implementing $G$ remains unchanged across iterations, the prior distribution $q_{0}$ also remains the same. However, the actual state distribution evolves with each iteration: after the $i$-th iteration, it is given by $p_{i}=Gp_{i-1}$, or more generally, $p_{i}=G^{i}p_{0}$. The total MMC then accumulates over iterations as:
@@
Even if the process starts at the prior $q_{0}$, after the first iteration, the distribution becomes $p_{1}=Gq_{0}$, which differs from $q_{0}$. This deviation from the prior distribution results in a strictly positive MMC in the next iteration, and the same holds for subsequent iterations wolpert2024thermodynamics .
Importantly, this strictly positive lower bound on EP holds for any initial distribution $p_{0}$, regardless of the prior $q_{0}$. As the system evolves through a sequence of distributions $\{p_{0},p_{1},\dots,p_{\tau}\}$, where $p_{t+1}=Gp_{t}$, there exists a distribution $\hat{q}_{0}$ that minimizes the expression on the right-hand side of Eq. 4. The associated MMC, denoted as $\mathcal{MC}_{\hat{q}_{0}}(p_{0})$, provides a strictly positive lower bound on the actual MMC, regardless of the prior distribution $q_{0}$.

### v3

@@ Abstract
All digital devices have components that implement Boolean functions, mapping that component’s input to its output. However, any fixed Boolean function can be implemented by an infinite number of circuits, all of which vary in their resource costs. This has given rise to the field of circuit complexity theory, which studies the minimal resource cost to implement a given Boolean function with any circuit. Traditionally, circuit complexity theory has focused on the resource costs of a circuit’s size (its number of gates) and depth (the longest path length from the circuit’s input to its output). In this paper, we extend circuit complexity theory by investigating the minimal thermodynamic cost of a circuit’s operation. We do this by using the mismatch cost (MMC) of a given circuit that is run multiple times in a row to calculate a lower bound on the entropy production (EP) incurred in each such run. Specifically, we discuss conditions under which MMC of a circuit is proportional to the size of a circuit, and when they are not. We also use our results to compare the thermodynamic costs of different circuit families implementing the same family of Boolean functions. In addition, we analyze how differences in the underlying physics of individual gates within a circuit influence the minimal thermodynamic cost of the circuit as a whole. These results lay the foundation for extending circuit complexity theory to include mismatch cost as a resource cost.
@@ I Introduction
Computational complexity theory explores the relative difficulty of computing functions within formal models of computation. In his foundational work [1], Alan Turing introduced the concept of Turing machines and demonstrated that functions can be classified as either computable or non-computable. Within the class of computable functions, it was later show that some are inherently harder to compute than others [2, 3]. Computational complexity theory [4] studies a fundamental question: what does it mean for one function to be more difficult to compute than another?
@@
A given function can often be computed using multiple algorithms, each incurring different resource costs depending on the model of computation. Traditionally, computational complexity theory has focused on quantifying resource usage such as the number of computational steps or memory requirements. Early discussions on complexity measures also considered physical energy expenditure and thermodynamic work as potential complexity metrics [5].
@@
Landauer and Bennett were among the first to explore the thermodynamic cost of computation using equilibrium statistical mechanics. They famously argued that erasing a single bit of information necessarily generates at least $k_{B}\log 2$ of heat, where $k_{B}$ is Boltzmann’s constant and $T$ is the temperature of the surrounding heat bath [6, 7]. However, this early work modeled computation—which is inherently a non-equilibrium process—within the framework of equilibrium thermodynamics. As a result, the analysis was semi-formal and limited in scope, leaving many fundamental questions unanswered. At the time, the theoretical tools needed to rigorously analyze far-from-equilibrium systems were not yet available, and the thermodynamics of computation remained an open and underdeveloped area.
@@
Fortunately, recent advances in non-equilibrium thermodynamics, particularly within the framework of stochastic thermodynamics, have made it possible to rigorously define and analyze thermodynamic quantities such as work, entropy, and heat dissipation in systems driven far from equilibrium [8, 9]. These tools are now being applied to fundamental questions about the thermodynamic costs of computation and communication processes [8, 9, 10, 11, 12]. In particular, stochastic thermodynamics allows us to analyze the entropy production, i.e., the dissipated heat, in a process.
@@
Boolean circuits provide an alternative model of computation, characterizing the complexity of a function in terms of the size and depth of the circuit implementing it. The size of a circuit is the total number of gates, while the depth corresponds to the longest path from an input to an output, representing the time required for computation [4, 13]. Just as there exist multiple algorithms for a given problem, there are infinitely many Boolean circuits that can implement a given Boolean function. The task of identifying the most efficient circuit among those infinitely many circuits is both theoretically difficult and practically important. Circuit complexity theory provides a framework for comparing and optimizing circuits based on size (number of gates) and depth (longest path from input to output), but the associated thermodynamic cost of running a circuit as a key measure in optimization has not been given considerable attention.
In this paper, we introduce the use of resource costs grounded in stochastic thermodynamics, in contrast to the traditional resource costs such as size and depth that have been analyzed for algorithms run on a Boolean circuit. Specifically, we focus on the use of mismatch cost (MMC) [14, 15, 16, 11, 17]. MMC is defined as the portion of total entropy production that arises when the actual distribution of inputs differs from the distribution a machine is optimized for, and it is largely independent of the microscopic physical details of the implementation. In this sense, MMC strengthens the second law by providing positive lower bounds on the energetic cost of running a computational machine. In the context of circuits, this implies positive lower bounds on the entropy production of running a Boolean circuit that hold regardless of the detailed physics of the gates, depending only on the circuit topology and the way computation proceeds through the gates, as we show in this paper.
For theoretical computer scientists, this work opens up a natural extension of circuit complexity by adding an energetic dimension based on MMC, thereby refining how one compares architectures that implement the same input–output map—not only by depth and size, but also by their associated MMC. Additionally, the scaling behavior of MMC with respect to problem size — analogous to other resource costs — opens up a particularly interesting direction. Moreover, other resource costs such as “time” and “space” that are of deep interest in CS theory are only defined for specific computational architectures, e.g., Turing Machines, Boolean circuits, or DFAs. While the definitions for these costs are analogous to one another, and in some cases are related up to polynomial transformations, they are formally different. In contrast, MMC is an architecture-independent resource cost: the exact same definition applies across computational architectures. Additionally, unlike purely abstract notions of computational complexity, MMC is partially informed by properties of the physical system implementing the computation, and can therefore be interpreted as a very coarse proxy for real-world energetic costs.
Circuit design provides an important example of the power of using MMC as the resource cost. A central challenge in digital circuit design is how to trade off different resource costs when implementing a desired computation [18, 19]. Energetic costs, and the associated heat dissipation, are a major concern. As processors grow smaller and more efficient [20] approaching thermodynamic limits, and as computational demands continue to grow–particularly due to large-scale data processing and machine learning workloads–understanding the energy cost of computation is becoming increasingly important [21]. Much prior circuit-level energy analysis relies on phenomenological power models and electrical dissipation estimates (e.g., CMOS power models and Joule heating in interconnects), rather than on thermodynamically grounded lower bounds [19, 22, 23].
@@
In contrast, incorporating non-equilibrium thermodynamic considerations into circuit complexity theory provides a principled framework for quantifying minimal heat dissipation in circuits. This perspective can refine existing complexity measures and help inform the design of architectures that are both logically and thermodynamically efficient. Moreover, since circuit structure determines whether a function is efficiently parallelizable or inherently sequential [24], thermodynamic analysis can shed light on the energy efficiency of parallel computation. Taken together, the framework we introduce provides a bridge between questions traditionally studied in statistical physics, computer science, and information theory, by placing them under a common energetic notion of computational cost.
@@ I.1 Contributions and roadmap
Sec. IV builds on this analysis to introduce mismatch cost complexity as a new circuit complexity measure, analogous to size and depth complexity, representing the minimal energetic cost of implementing a Boolean function with any circuit. We derive results and theorems that link MMC complexity to size complexity, showing that under certain conditions, the upper bound on MMC grows linearly with the size complexity of a circuit family. Complementing this result, we identify conditions where MMC scales differently from size complexity. In Sec.V, we discuss the energetic aspects of circuit optimization by comparing the MMC associated with circuit families that implement a given Boolean function. Specifically, we study two families that compute the ADD function—one with depth complexity $\log(n)$ and another with $n$. We also investigate the impact of heterogeneity in the physical properties of gates on the associated MMC. Finally, Sec. VI concludes the paper by discussing possible directions for future research.
I.2 Earlier work
The tools of stochastic thermodynamics have been applied to computational models such as finite automata [11, 25, 16] and Turing machines [26, 27], providing one of the first principled frameworks for analyzing the thermodynamic costs of computation. Specific to circuits, prior work has estimated the heat dissipated during gate operations using Landauer’s bound [28]. While this was a pioneering approach, stochastic thermodynamics now recognizes that the Landauer bound—which accounts only for entropy changes—can be zero even when the total dissipated heat is substantial [29]. This highlights the importance of entropy production (EP), a central concept in stochastic thermodynamics.
More recent studies have incorporated EP into the analysis of Boolean circuits by modeling gate dynamics as continuous-time Markov chains (CTMCs) [30]. However, such a framework requires properly incorporating dynamics over hidden states, which in general are necessary for there to be a CTMC that can implement discrete-time dynamics [31, 31]. In contrast, our analysis does not assume any specific form for the underlying continuous-time dynamics. Furthermore, in our model, the logical values at gates are not reset between successive runs, more closely reflecting the behavior of many physical computing devices. In contrast, prior work [30] assumes a full reset after each computation.
Another closely related set of results analyzes entropy production in stochastic processes structured as Bayesian networks [32, 33, 34]. These works provide general decompositions of entropy production in terms of information flows along the edges of a directed acyclic graph. The formalism developed in those works is in principle extensible to circuit-like architectures, and our framework can be viewed as a specialization of these more general thermodynamic decompositions to Boolean circuits, together with a complexity-theoretic perspective on how costs scale with circuit size.
@@ II.1 Circuit complexity theory
In this section, we provide a brief and necessarily compact overview of basic concepts from circuit complexity theory. For a more detailed and comprehensive treatment, see [4].
A circuit is built from basic logic gates that are interconnected to process inputs and produce outputs (Fig. 1). A basis is a set $\mathcal{B}$ of Boolean functions that define the allowable gates in a circuit. For example, A basis containing AND and NOT gates is considered a universal basis because any Boolean function can be implemented using only these gates. Formally, a Boolean circuit is represented as a directed acyclic graph (DAG), denoted by $(V,E,\mathcal{B},\mathcal{X})$, where $V$ is the set of nodes representing a logic gate selected from a finite basis $\mathcal{B}$. The set of edges $E$ define the directed connections between gates, determining how information flows through the circuit. The set $\mathcal{X}$ represents the state space of the circuit (see below). The direction of edges reflects dependencies between gates, enabling a specific ordering of nodes in $V$. This ordering, called a topological ordering, ensures that for every directed edge from node $\mu$ to node $\nu$, node $\mu$ appears before node $\nu$ in the sequence ($\mu<\nu$).
A node $\mu$ has incoming edges from nodes known as its parent nodes, denoted as $\mathrm{pa}(\mu)$, and outgoing edges to nodes known as its children nodes, denoted as $\mathrm{ch}(\mu)$. Nodes that have no incoming edges, called input nodes or root nodes, serve as the circuit’s inputs. We denote the set of input nodes as $V_{\mathrm{in}}\subset V$ and the set of all non-input nodes as $V_{\mathrm{nin}}=V\setminus V_{\mathrm{in}}$. Each gate $\mu$ in the circuit has an associated state space, denoted by $X_{\mu}$, and its state is represented by $\boldsymbol{x}_{\mu}\in X_{\mu}$. For binary gates, the state space is $X_{\mu}=\{0,1\}$. The collective states of all gates together define the joint state of the circuit. The joint state space of the circuit is given by
@@
We use $\boldsymbol{x}_{\mathrm{pa}(\mu)}$ and $\boldsymbol{x}_{\mathrm{ch}(\mu)}$ to denote the joint state of the parent and children gates of a gate $\mu$, respectively. Similarly, we define $\boldsymbol{x}_{\mathrm{in}}$ and $\boldsymbol{x}_{\mathrm{nin}}$ to represent the joint states of all input nodes and all non-input nodes, respectively.
@@
Circuits differ fundamentally from other models of computation. In models like Turing machines, a single machine accepts input of of any arbitrary length. In contrast, circuits require a distinct specification for each input length. This means that to compute functions belonging to the same family but with different input sizes, one must design distinct circuits for each case. Because the model does not provide a uniform mechanism for handling all input lengths, it is classified as a non-uniform model of computation [4, 13]. As a result, computation in this model is described using a circuit family, denoted as $\{C_{n}\}_{n\in\mathbb{N}}$, where each circuit $C_{n}$ corresponds to inputs of length $n$. A circuit family $\mathcal{C}$~ is a sequence $\{C_{n}\}_{n\in\mathbb{N}}$ of Boolean circuits, where $C_{n}$ is the circuit with $n$ inputs.
@@
Analogously, let $d:\mathbb{N}\to\mathbb{N}$. 0ptd0pt{d} is the class of functions $f$ for which there exist a circuit family $\mathcal{C}$~ of depth complexity $d$.
The growth of a circuit’s size or depth with respect to the input length of a function provides a crucial measure of the function’s computational complexity. In his seminal work, Shannon proved that most Boolean functions require circuits of exponential size, specifically showing that there exist functions that cannot be computed by circuits smaller than $2^{n}/10n$. In other words, the vast majority of functions—often referred to as ”hard” functions—such that $f\notin\mathbf{SIZE}(2^{n}/10n)$ and thus require exponentially large circuits to compute [35]. However, despite this result, explicitly identifying such a function has remained challenging, as most functions of practical interest can still be computed by circuits of reasonable, albeit large, size.
An interesting class of functions emerges when restricted to ”small” circuits, i.e., circuit families whose size does not grow faster than polynomial with input length. $\mathbf{P_{/poly}}$ is the class of functions for which there exist a $\mathrm{poly}(n)$-size circuit family,
@@
A function has an efficiently parallel algorithm if and only if it is in $\mathbf{NC}$ [4, 13].
The previous example of ADD is in $\mathbf{NC}$. However, just like any Boolean function, ADD has many circuit families that can implement it. For example, ADD can be implemented using two well-known circuit families: the ripple-carry adder (RCA) and the carry look-ahead adder (CLA), each offering distinct trade-offs in size and depth complexity [13].
The RCA follows the simple ”carry-over” method taught in high school arithmetic, where each bit addition depends on the carry from the previous bit. As a result, for an $n$-bit input, the depth of the circuit grows linearly with $n$. Additionally, increasing the input size by one bit requires adding one additional full-adder circuit to the chain. This means that the overall circuit size also grows linearly with $n$. Its simplicity makes it easy to scale. However, since each stage of bit addition must wait for the carry from the previous stage before proceeding, RCA becomes significantly slower as the input size increases.
In contrast, CLA uses a more sophisticated approach by predicting carry bits in advance rather than waiting for them to propagate sequentially . This significantly reduces the depth of the circuit, which grows proportionally to the logarithm of the input size. As a result, computation time decreases drastically compared to RCA. However, the parallel computation and complex wiring cause the total number of gates in the circuit to grow at a rate proportional to $\log n$. Consequently, as the input size increases, the CLA circuit becomes increasingly intricate.
@@ II.2 Stochastic Thermodynamics and Mismatch Cost
Real-world computational systems operate far from equilibrium, with multiple interdependent components evolving rapidly. Traditional equilibrium statistical mechanics, which applies only to large macroscopic systems that either evolve quasi-statically or remain static, is inadequate for analyzing the energetic costs of computation [12]. Mesoscopic systems far from equilibrium that exhibit stochastic fluctuations can be analyzed within the framework of stochastic thermodynamics. The key assumption in this framework is that any degrees of freedom not explicitly described by the system’s dynamics—such as internal states or the reservoirs—remain in equilibrium. This assumption allows one to relate thermodynamic quantities to system dynamics through the principle of local detailed balance [9, 36, 37, 38].
One of the central quantities in stochastic thermodynamics is entropy production, which quantifies the irreversible heat dissipation in a process. If the probability distribution over the states of a physical system during its evolution changes from $p_{0}$ to $p_{\tau}$, the total EP associated with an initial distribution $p_{0}$ can be expressed as [8, 15],
@@
Another important contribution to the EP of running a system is its “mismatch cost”. Consider the initial distribution $q_{0}$ that minimizes the EP generated by running a system, given in Eq. (1). In the literature, this distribution is referred as the “prior” distribution of the running the system. The reason for this terminology is that $q_{0}$ plays the role of a Bayesian prior for an agent trying to change the distribution that the system has at the end of running the system back to the original distribution that the system had before the system ran — all while expending the minimal amount of free energy. In particular, that agent must perform a Bayesian inverse to estimate what that original distribution over the states of that system were before that system ran, and $q$ is the proper choice for the prior of this Bayesian inverse [39].
When the process starts from a different initial distribution $p_{0}$, the resulting EP, $\sigma(p_{0})$, can be decomposed as [15, 14],
@@
where $D(p_{0}||q_{0})$ is the Kullback-Leibler (KL) divergence between $p_{0}$ and $q_{0}$, and $\sigma_{\text{res}}$, known as the residual EP, represents the minimum EP of the process, defined as $\sigma_{\text{res}}:=\sigma(q_{0})$. The drop in KL-divergence $D(p_{0}||q_{0})-D(Gp_{0}||Gq_{0})$ is called the mismatch cost. The MMC is always non-negative, a consequence of the data processing inequality for KL-divergence. Unlike the residual EP $\sigma_{\text{res}}$, which depends on the fine details of the physical implementation, MMC depends only on the computational map $G$, the initial distribution $p_{0}$, and the prior distribution $q_{0}$. All of the physical details of the system are encoded in the choice of the prior. Additionally, $\sigma_{\text{res}}$ is non-negative due to the second law. In certain dynamical situations, the MMC provides a strictly positive contribution to EP.
Consider a process where the computational map $G$ is repeatedly applied to the state space. Since the underlying physical process implementing $G$ remains unchanged across iterations, the prior distribution $q_{0}$ also remains the same. However, the actual state distribution evolves with each iteration: after the $i$-th iteration, it is given by $p_{i}=Gp_{i-1}$, or more generally, $p_{i}=G^{i}p_{0}$. The total MMC then accumulates over iterations as:
@@
Even if the process starts at the prior $q_{0}$, after the first iteration, the distribution becomes $p_{1}=Gq_{0}$, which differs from $q_{0}$. This deviation from the prior distribution results in a strictly positive MMC in the next iteration, and the same holds for subsequent iterations [11].
Importantly, this strictly positive lower bound on EP holds for any initial distribution $p_{0}$, regardless of the prior $q_{0}$. As the system evolves through a sequence of distributions $\{p_{0},p_{1},\dots,p_{\tau}\}$, where $p_{t+1}=Gp_{t}$, there exists a distribution $\hat{q}_{0}$ that minimizes the expression on the right-hand side of Eq. (4). The associated MMC, denoted as $\mathcal{MC}_{\hat{q}_{0}}(p_{0})$, provides a strictly positive lower bound on the actual MMC, regardless of the prior distribution $q_{0}$.
@@
Figure 3: Evolution of joint distribution under re-initialization of inputs and re-running the circuit gate-by-gate. (A) After a complete run, the joint distribution reflects maximal correlation among all gates. This correlation arises from the initial input distribution and the network of dependencies within the circuit. (B) In the next run, the input nodes are re-sampled from the input distribution $p(\boldsymbol{x}_{\mathrm{in}})$, making the new inputs independent of the states of non-input gates from the previous run. As a result, the joint distribution after re-initialization is factored as $p(\boldsymbol{x}_{\mathrm{in}})p(x_{1},...,x_{5})$, indicating that inputs and non-input gates are now statistically independent. (C) When gate $x_{1}$ updates based on the newly re-sampled input values, it becomes correlated with the input nodes while simultaneously losing correlation with the rest of the gates. Consequently, the joint distribution evolves to $p(\boldsymbol{x}_{\mathrm{in}},x_{1})p(x_{2},...,x_{5})$. Similarly, (D), (E), (F), and (G) illustrate the sequential updates of the remaining gates and the corresponding evolution of the distribution. In particular, (G) demonstrates that after a complete run of the circuit, the distribution returns to its initial form as shown in (A).